OpenStudy (anonymous):

solve (t^2+6t)/t-1 - (7)/t-1

OpenStudy (anonymous):

factor out the numerator.. multiply both expressions by the LCD which is t-1 so you can not deal with fractions.. then solve for x.

OpenStudy (anonymous):

\[\frac{ t ^{2}+6t }{ t-1 }-\frac{ 7 }{ t-1 }\]Is that the correct equation

OpenStudy (anonymous):

er t

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

I would not do what nameless said. it over complicates the problem

OpenStudy (anonymous):

since the both have the same denominaor already you can add the numerators

OpenStudy (anonymous):

in this case subtract them

OpenStudy (anonymous):

\[\frac{ t ^{2}+6t-7 }{ t-1 }\]

OpenStudy (anonymous):

now when we factor we need to think about was numbers, when multiplied equal -7. we get -7, 1 and -1, 7.

OpenStudy (anonymous):

then we look at these numbers and think what numbers when added will give us 6.

OpenStudy (anonymous):

we should arrive at \[\frac{ (t+7)(t-1) }{ t-1 }\]

OpenStudy (anonymous):

as you can see you can cancel off the t-1's and your left with t+7

OpenStudy (anonymous):

Ahh that is a nice way to do that. Thanks for helping me as well :)

OpenStudy (anonymous):

Do you understand? ur welcome @Nameless

OpenStudy (anonymous):

yes i understand what you saying